OpenStudy (anonymous):

–2x + 3y = -6 can someone please put this in slope intercept form for me. I need help. I keep getting xy

OpenStudy (campbell_st):

start by adding 2x to both sides, .... what do you get..?

OpenStudy (anonymous):

3y=2x-6 y=(2x-6)/3 y=2/3x-6/3 y=2/3x-2

OpenStudy (anonymous):

I got 3y= -4x

OpenStudy (campbell_st):

ok... lets start again, do it this way, add 2 x to both sides -2x + 3y = -6 +2x +2x ----------------- 3y = 2x - 6 notice that the 2x and 6 are not alike... so that can't be added or subtracted. the last step is to divide both sides by 3 can you do that..?

OpenStudy (anonymous):

I got y=2x-2.......

OpenStudy (campbell_st):

nearly when you divide 2 by 3 you get 2/3 so the solution is y = 2/3 x - 2 the slope is 2/3 and the y-intercept is y = -2 hope it helps

OpenStudy (anonymous):

oh i get it now thank you